Circular No. 7752

Central Bureau for Astronomical Telegrams INTERNATIONAL ASTRONOMICAL UNION

Mailstop 18, Smithsonian Astrophysical Observatory, Cambridge, MA 02138, U.S.A.

IAUSUBS@CFA.HARVARD.EDU or FAX 617-495-7231 (subscriptions) CBAT@CFA.HARVARD.EDU (science)

        URL http://cfa-www.harvard.edu/iau/cbat.html ISSN 0081-0304

Phone 617-495-7440/7244/7444 (for emergency use only)

SUPERNOVAE 2001fc, 2001fd, 2001fw, AND 2001fy A. V. Filippenko and R. Chornock, University of California at Berke-

ley, write: “Inspection of CCD spectra (range 330–1000 nm), obtained on Nov. 16 UT with the Shane 3-m reflector at Lick Observatory, reveals that SN 2001fc (IAUC 7741, 7743) is a young type-II supernova; weak, broad Hα emission is visible, but the spectrum is generally featureless. The red continuum suggests considerable extinction; on the other hand, the interstellar Na I D lines are relatively weak, and the object is not unusually dim, given its NASA/IPAC Extragalactic Database (NED) redshift of 5000 km/s. It is possible that SN 2001fc is a type-IIb supernova like SN 1987K (Filippenko 1988, A.J. 96, 1941) at an early stage of development. The spectrum of SN 2001fw (IAUC 7751) also resembles that of SN 1987K, but with very weak Hα emission; its exact classification (IIb, Ib, or Ic) is uncertain. SN 2001fd (IAUC 7742) is of type II, with strong, broad Hα emission; Hβ and Hγ exhibit P-Cyg profiles. The spectrum of SN 2001fy (IAUC 7751), also of type II, looks very similar to that of SN 2001fd but with a bluer continuum.”

SUPERNOVA 2001fe IN UGC 5129 T. Matheson, S. Jha, P. Challis, and R. Kirshner, Harvard-Smithsonian

Center for Astrophysics, report that a spectrum (range 370–750 nm) of SN 2001fe (cf. IAUC 7742), obtained by P. Berlind on Nov. 14.49 UT with the F. L. Whipple Observatory 1.5-m telescope (+ FAST spectrograph), shows it to be a type-Ia supernova near maximum light. Adopting the NED recession velocity of 4059 km/s for the host galaxy, the supernova expansion velocity is ∼ 11 200 km/s for Si II (rest 635.5 nm). The spectralfeature age of the supernova (Riess et al. 1997, A.J. 114, 722) is 1 ± 2 days before maximum light.

Filippenko and Chornock also report that inspection of CCD spectra, obtained at Lick as above on Nov. 15 UT, reveals that SN 2001fe is of type Ia, close to maximum brightness.

SUPERNOVA 2001ff IN UGC 4685 Matheson et al. add that a spectrum (range 370–750 nm) of SN 2001ff

(cf. IAUC 7743), obtained by Berlind as above on Nov. 14.47 UT, shows it to be a type-II supernova past maximum light. Adopting the NED recession velocity of 3978 km/s for the host galaxy, the expansion velocity derived from the minimum of the Hα line is 8400 km/s.

2001 November 16  \begin{matrix}c \\ ◯ \end{matrix} Copyright 2001 CBAT  Daniel W. E. Green